OpenStudy (anonymous):

Ben wishes to strength a mixture from 10% alcohol to 30% alcohol. How much pure alcohol should be added to 7L of the 10% mixture?

OpenStudy (anonymous):

let x = the amount of alcohol added, youll get this equation: (0.7 + x) / (7+x) = 30/100, 70 + 100x = 210 + 30x 70x = 140 x=2

OpenStudy (anonymous):

This guy is correct ^^